# Multi-scale Laplacian community detection in heterogeneous networks

### v3

@@ Abstract
Heterogeneous and complex networks represent the intertwined interactions between real-world elements or agents. A fundamental problem of complex network theory involves finding inherent partitions, clusters, or communities. By taking advantage of the recent Laplacian Renormalization Group approach, we scrutinize information diffusion pathways throughout networks to shed further light on this issue. Based on inter-node communicability, our definition provides a unifying framework for multiple partitioning measures: the multi-scale Laplacian (MSL) community detection algorithm. This new framework, free of topological null-model assumptions, permits to introduce a scale-dependent optimal partition in communities and to determine the existence of a particular class of nodes, called metastable nodes, that switching communities at different scales, are expected to play a central role in the communication between different communities and, therefore, in managing macroscopic effects of the whole network.
@@
Detecting mesoscale structures (i.e., communities) in networks has become a fundamental problem –and still a matter of debate [24]– in network science, where many different methods have been proposed and classified according to their performance [25]. However, the problem lacks a holistic perspective, i.e., a unified theoretical framework capable of setting the limitations of the different methods. For instance, modularity has been demonstrated to be a linear approximation of Markov stability [26]. Still, community detection algorithms based on quality functions strongly depend on the null models they must consider to meaningfully quantify their statistical significance [26, 27]. Another crucial problem lies in the fact that the definition of community in general is strictly connected to functional properties and, thus, to the process on the network for which we want to find the optimal partition [28]. Moreover, the proper identification of local communities represents a fundamental open problem when tackled through usual community detection methods [24]. In a nutshell, identifying basic structural and dynamic (functional) communities in a rigorous and universal way, without any topological null-model assumption, raises a general problem of definition that still needs to be resolved.
@@
Here, in the light of the LRG, we develop a new community detection framework that unravels the complete hierarchical structures of the network at different resolution levels. In particular, we reconstruct the hierarchical network dendrogram of clusters based on communication distances among the nodes by taking advantage of the LRG in their $k-space$ formulation without needing a priori null models or making statistical inference. Therefore, we propose the optimal partition as the one determined by the slowing down of the information diffusion throughout the network. In particular, we also show the existence of metastable nodes that can be members of different communities at different intrinsic network timescales. Finally, we illustrate the emergence of functional communities when a set of random walks explore the network structure and discuss the bridges with previous proposals, unifying modularity maximization, Markov stability, and spectral methods in a common framework.
I Multi-scale Laplacian community detection in the framework of the Laplacian Renormalization Group
Here we propose a new framework of community detection, the multi-scale Laplacian (MSL), which can correctly detect communities also in challenging cases and shares with the Markov Stability the property of being time-dependent and defined through the evolution operator of a diffusion process. However, in our case, we ground our approach on the symmetric Laplacian, $\hat{L}=\hat{D}-\hat{A}$, where $\hat{D}$ is the degree matrix and ${{\hat{A}}}$ the adjacency matrix of the network. Regarding the graph Laplace/heat equation,
$\dot{\textbf{s}}(\tau)=-\hat{L}\textbf{s}(\tau)\,,$
we have specifically considered its time evolution operator $e^{-\tau\hat{L}}$, which is on the basis of a canonical description of heterogeneous networks.
In particular, as well shown in two recent papers [33, 17], the Laplacian density matrix, initially introduced in [34, 35]
@@ I Multi-scale Laplacian community detection in the framework of the Laplacian Renormalization Group
permits through the related Laplacian entropy
$S(\tau)=-\frac{1}{\log N}\mbox{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]=-\frac{\left<\log\rho(\tau)\right>_{\tau}}{\log N}$ (2)
and the entropic susceptibility (or specific heat)
$C(\tau)=-\frac{dS}{d\log\tau}\,,$
to accurately study the multi-scale organization of any network. Indeed, these quantities are able to detect the information core of the network, the related structural/diffusional phase transitions [33], and to naturally extend the statistical physics coarse-graining and Renormalization Group to heterogeneous networks through the formulation of the LRG [17] (see also Appendix A).
@@
Figure 1: Sketch of the MSL framework. Communication distance between nodes reflects the underlying hidden complex topology, capturing the inter-nodes communicability and giving a natural merging of nodes at specific times, $\tau$. The LRG gives a natural interpretation of the network scales in terms of network eigenmodes, thereby enabling a deeper understanding of the mesoscopic properties of the network.
Note that the operator $e^{-\tau\hat{L}}$ accounts for the sum of diffusion trajectories along all possible paths connecting nodes $i$ and $j$ in a temporal scale $\tau$ [22, 36]. The operator $\hat{K}(\tau)=e^{-\tau\hat{L}}$ or, equivalently $\hat{\rho}(\tau)$, gives a robust measure of information communicability or diffusion strength between pair of nodes on a time scale $\tau$. Consequently, using this canonical ensemble formulation of a heterogeneous topology, we can: (i) define at each time $\tau$ an ultrametric distance between nodes from the operator $\hat{K}(\tau)$, (ii) build a clustering dendrogram, (iii) introduce a consistent criterion for an optimal cut of the dendrogram from which we can derive the best partition at time $\tau$, and finally (iv) study how this changes by varying $\tau$.
@@
The first step consists of defining a Communication distance at time $\tau$. From the physical meaning of $e^{-\tau\hat{L}}$ it is natural to define such symmetric distance between any pair of nodes $i$ and $j$ as $\mathcal{D}_{ij}(\tau)=\frac{1-\delta_{ij}}{K_{ij}(\tau)}$ (as shown below, this can be equivalently defined as $\mathcal{D}_{ij}(\tau)=\frac{1-\delta_{ij}}{\rho_{ij}(\tau)}$). Indeed $K_{ij}(\tau)$ gives the fraction of initial information on node $i$ that is transferred to node $j$ by diffusion in a time $\tau$ and vice-versa. Equivalently, one can write $K_{ij}(\tau)=\frac{\delta s_{j}(\tau)}{\delta s_{i}(0)}$, where $\delta s_{j}(\tau)$ is the fluctuation of the information on node $j$ at time $t$ induced by a fluctuation of the initial information $\delta s_{i}(0)$ on node $i$, keeping as fixed the initial condition on the other nodes. Consequently, we can say that the larger $\mathcal{D}_{ij}(\tau)$, the less the communicability on a time $\tau$ between nodes $i$ and $j$.
Note that $\mathcal{D}_{ij}(\tau)$ defines an ultrametric distance, as it satisfies the following conditions at all $\tau>0$: (i) $\mathcal{D}_{ij}\geq 0$ with $\mathcal{D}_{ij}=0$ if $i=j$; (ii) $\mathcal{D}_{ij}=\mathcal{D}_{ji}$, and (iii) $\mathcal{D}_{ij}\leq\max\left(\mathcal{D}_{iz},\mathcal{D}_{zj}\right)$ $\forall z\in{\cal V}$ (where ${\cal V}$ is the set of nodes of the network).
@@
Once we have introduced the distance $\mathcal{D}_{ij}(\tau)$ between any pair of nodes at each time $\tau>0$, we can formulate a simple method for grouping nodes into clusters at arbitrary $\tau$. We emphasize that, even though complex networks lack an explicit spatial embedding, $\mathcal{\hat{D}}$ allows us to reduce the problem to the traditional clustering problems in a metric space [37] through the construction of a dendrogram  [^1] In fact, $\mathcal{\hat{D}}$ is a projection of how points (nodes) are distributed in the N-dimensional k-space representation of the network.. In order to define a stable algorithm for the construction of the dendrogram, we choose the average group clustering algorithm, a compromise between the sensitivity to outliers of the complete clustering and the tendency of single-linkage clustering to inhibit compact clusters [39, 37]. In this way, the merging of two clusters happens when the mean distance $\overline{\mathcal{D}}$ between any node belonging to the first cluster and any node belonging to the second cluster becomes smaller than the chosen threshold $\Delta$. The output of this algorithm will thus be a hierarchical tree for each resolution scale $\tau$, without any null-model assumption. At this point, three natural questions arise: (1) Given the dendrogram at diffusion time $\tau$, what is the best partition, i.e., the best “cut” of the hierarchical clustering tree? (2) Since the metric $\mathcal{\hat{D}}(\tau)$ may, at different $\tau$, determine different dendrograms in which the specific length of branches and their composition can change (other metrics shares this situation; e.g. the same applies to Markov stability [21]), is there an optimal choice of $\tau$? (3) What is the stability of the optimal cut, i.e. optimal partition, at varying $\tau$? In this section, we face the first two questions and leave the third one to Sect. III.
@@
Figure 2(a) shows the dendrogram resulting from our method in the case of Zachary’s karate club [2, 40], which serves as a benchmark showing the usual partition of the network in two communities. Figure 2(b) shows the entropic phase transition [33] for Zachary’s karate club through the behavior of the Laplacian entropy $S(\tau)$ in time (red dashed line) and of the entropic susceptibility (i.e., the specific heat) $C(\tau)$.
@@
Let us now face the first question above, which consists of quantifying the “quality” of the different network partitions encrypted in a single dendrogram (i.e., dendrograms cut at different levels of $\Delta$) at a fixed diffusion time $\tau$. We formulate a comparative criterion to select the “optimal” partition by introducing a suitable Partition Stability Index $\Psi(n;\tau)$. It is defined as the ratio between the dendrogram gaps and the total dendrogram length in log scale:
@@
where $\Delta_{n}(\tau)$ is the value of the threshold $\Delta$ corresponding to the $n^{th}$ dendrogram branching, starting from the network as a single cluster (note that with this choice of the index $n$, $\Delta_{n}(\tau)$ is a decreasing function of $n$) and $\mathcal{N}=[\log\Delta_{1}(\tau)-\log\Delta_{n_{max}}(\tau)]^{-1}$ is a normalization constant. The quantity $\Psi(n;\tau)$ measures the minimal distance (dissimilarity) in logarithmic scale between all the pairs of clusters respectively belonging to level $n+1$ and to level $n$. The use of the log scale is motivated by the definition of the entropy $S(\tau)\sim\left<\log\hat{\rho}(\tau)\right>_{\tau}$ itself: the higher $\Psi(n;\tau)$, the higher the (Laplacian) information gap between the two consecutive branching of indices $n$ and $n+1$. It is, therefore, natural to consider the optimal cut of the dendrogram at the value $n$ maximizing $\Psi(n;\tau)$. Indeed, the network partition at $n$ maximizing $\Psi(n;\tau)$ contains the maximal Laplacian information induced by the topology of the network. In the case of multiple intrinsic scales of the network, $\Psi(n;\tau)$ can present different local maxima. All of them indicate significant partitions of the network, the absolute maximum being the one to be compared with other standard community detection algorithms.
@@
Figure 2: MSL communities. (a) Normalized dendrogram for the Zachary’s karate club with $\tau\lesssim\tau^{\prime}=1/\lambda_{max}$. Red dashed line reflects the optimal division as stated by the maximum of $\Psi$. Different communities are shaded in different colors. (b) Entropy parameter (red dashed line, $(1-S)$), and specific heat (blue solid line, $C$), versus the temporal resolution parameter of the network, $\tau$. Black dashed line indicates $\tau^{\prime}=1/\lambda_{\max}$. Black dotted line indicates the time of the maximum heat capacity, $(\tau^{*})$. (c) Partition Stability Index $(\Psi)$ versus gap number for $\tau\lesssim\tau^{\prime}$. Note the high $\Psi$ values giving place to the usual division in two and four communities of Zachary’s karate club. Inset: Optimal division into communities of the network.
Once we have determined how to locate the “best cut” of the dendrogram at fixed $\tau$, we move to face the problem of comparing dendrograms and relative cuts at different $\tau$’s.
@@
This can be done by exploiting the recently introduced theory of LRG [17]. It grounds on the fact that the network Laplacian $\hat{L}$ –and in particular its spectrum of eigenvalues $\lambda$ and eigenvectors |$\ket{\lambda}$– encodes the whole information about all the independent network diffusion modes (see Appendix B and Appendix C for further clarifications). As explicitly shown in [33] and [17], these modes collected in the operator $e^{-\tau\hat{L}}$ determine the clusters of nodes that are informationally connected at the time scale $\tau$. Indeed, the LRG scheme takes advantage of the fact that, by varying the diffusion time $\tau$, the operator $\hat{\rho}(\tau)$ works as a sort of “scanner” of the characteristic scales of the network [41]. Such mesoscopic scales can be detected by progressively integrating out smaller and smaller Laplacian eigenmodes and, consequently, the corresponding network structures as $\tau$ increases. For finite but large networks, this procedure, and in particular the susceptibility $C(\tau)$, generally determines the existence of a resolution window where the different partitions of the network in clusters are relevant. The smallest possible network scale corresponds to the maximum eigenvalue $\lambda_{max}$: the choice of a diffusion time-scale $\tau\sim\nicefrac{{1}}{{\lambda_{max}}}$ coincides with the finest possible resolution of the network structures (this sheds light on the previously enigmatic origin of the resolution limit [42], see below). Starting from this scale $(\tau^{\prime}=1/\lambda_{max})$, one can consider progressively larger diffusion times integrating out smaller network eigenmodes up to $\tau\sim 1/\lambda_{gap}$, where $\lambda_{gap}$ is the smallest positive eigenvalue, also called Fiedler eigenvalue or gap, and monitoring the stability across diffusion scales of the partition of the network into clusters (i.e., defocusing microscopic details). This is related to the maximal characteristic network scale marked by the specific heat peak at long times $(\tau^{*})$, beyond which it rapidly decreases to zero.
@@
Figure 2(c) shows the behavior of $\Psi(n)$ for Zachary’s karate club and the inset the optimal partition obtained by our method for the time $\tau$ indicated by the vertical dashed line in Fig. 2(b).
A paradigmatic application of the method is the study of the structural organization of heterogeneous hierarchical nested networks with multiscale communities. In this context, we analyze both hierarchical-modular networks with a preferential attachment rule (HM-CP, that produces a core-periphery structure involving central connector hubs having local and global rich clubs [48]) and hierarchical lattices as the Dorogovtsev-Goltsev-Mendes graph (DGM), a pseudo-fractal graph with high clustering properties [49]. As better illustrated in Appendix D, our LRG-based approach can capture the full intrinsic modular structure of nested hierarchical networks (HM-CP), even when the inter-module connections and communication paths tend to be centralized through the hubs, as observed in real neural and brain networks [50]. The same happens also for the highly hierarchical DGM graph, whereas usual community detection methods present severe issues in unraveling either communities and the complete hierarchical structure of the network (see Appendix D). Figures 3(a) and (b) illustrate the emergent nested structure already apparent in the distance matrix, $\mathcal{\hat{D}}$, for HM-CP and DGM networks. Finally, we have successfully tested the Lancichinetti-Fortunato-Radicchi benchmark [51], a network generator with a priori known heterogeneous communities, which serves as a stringent criterion, to compare different community detection algorithms (see Appendix D for further details).

### v4

@@ Abstract
Heterogeneous and complex networks represent intertwined interactions between real-world elements or agents. Determining the multi-scale mesoscopic organization of clusters and intertwined structures is still a fundamental and open problem of complex network theory. By taking advantage of the recent Laplacian Renormalization Group, we scrutinize information diffusion pathways throughout networks to shed further light on this issue. Based on inter-node communicability, our definition provides a clear-cut framework for resolving the multi-scale mesh of structures in complex networks, disentangling their intrinsic arboreal architecture. As it does not consider any topological null-model assumption, the LRG naturally permits the introduction of scale-dependent optimal partitions. Moreover, we demonstrate the existence of a particular class of nodes, called metastable nodes, that switching regions to which they belong at different scales, likely playing a pivotal role in cross-regional communication and, therefore, in managing macroscopic effects of the whole network.
@@
Detecting mesoscale structures (i.e., communities) in networks has become a fundamental problem –and still a matter of debate [24]– in network science, where many different methods have been proposed and classified according to their performance [25]. However, the problem lacks a holistic perspective, i.e., a unified theoretical framework capable of setting the limitations of the different methods. For instance, modularity has been demonstrated to be a linear approximation of Markov stability [26]. Besides, community detection algorithms based on quality functions strongly depend on the null models they must consider to meaningfully quantify their statistical significance [26, 27]. Another crucial problem is that the definition of a community in general is strictly connected to functional properties and, thus, to the process on the network for which we want to find the optimal partition [28]. Moreover, the proper identification of local communities represents a fundamental open problem when tackled through usual community detection methods [24]. In a nutshell, identifying basic structural and dynamic (functional) communities disentangling the arboreal architecture of a complex network without any topological null-model assumption, raises a general problem of definition that still needs to be resolved.
@@
Here, grounding on the LRG, we unravel the entire multi-scale structure of every network at different resolution levels. In particular, we reconstruct the hierarchical network dendrogram (the arboreal architecture of the network), proposing a new concept of communication distances among nodes by taking advantage of the LRG in their $k-space$ formulation without needing a priori null models or statistical inference. Therefore, we show how the optimal partitions are those determined by the slowing down of the information diffusion throughout the network. In particular, we evidence the existence of metastable nodes that can dynamically switch the region to which they belong at different intrinsic network timescales. Finally, we illustrate the emergence of functional dendrograms when a set of random walks explore the network structure and discuss the bridges with previous frameworks, such as modularity maximization, Markov stability, and spectral methods, opening new avenues in future developments in the networks field.
I Real-Space distances of non-euclidean topological structures
We ground our approach on the recently proposed LRG [33, 17], and in the Laplacian density matrix [34, 35],
@@ I Real-Space distances of non-euclidean topological structures
where $\hat{L}=\hat{D}-\hat{A}$ represents the symmetric Laplacian, $\hat{D}$ is the degree matrix and $\hat{A}$ the adjacency matrix of the network. In particular, the entropic susceptibility (or specific heat, see Appendix A) $C(\tau)=-\frac{dS}{d\log\tau}$, allows us to analyze the multi-scale organization of any network accurately [33, 17]. The operator $\hat{K}(\tau)=e^{-\tau\hat{L}}$ or, equivalently $\hat{\rho}(\tau)$, gives a robust measure of information communicability or diffusion strength between pair of nodes on a time scale $\tau$, accounting for the sum of diffusion trajectories along all possible paths connecting nodes $i$ and $j$ in a temporal scale $\tau$ [22, 36].
@@
Figure 1: Defining multi-scale Kadanoff supernodes. Communication distance between nodes reflects the underlying hidden complex topology, capturing the inter-nodes communicability and giving a natural merging of nodes at specific times, $\tau$. The LRG gives a natural interpretation of the network scales in terms of network eigenmodes, thereby enabling a deeper understanding of the mesoscopic properties of the network.
@@
First, we propose a Communication distance at time $\tau$. From the physical meaning of $e^{-\tau\hat{L}}$ it is natural to define such symmetric distance between any pair of nodes $i$ and $j$ as $\mathcal{D}_{ij}(\tau)=\frac{1-\delta_{ij}}{K_{ij}(\tau)}$ (or equivalently $\mathcal{D}_{ij}(\tau)=\frac{1-\delta_{ij}}{\rho_{ij}(\tau)}$). Indeed, $K_{ij}(\tau)$ gives the fraction of initial information on node $i$ that is transferred to node $j$ by diffusion in a time $\tau$ and vice-versa. Equivalently, one can write $K_{ij}(\tau)=\frac{\delta s_{j}(\tau)}{\delta s_{i}(0)}$, where $\delta s_{j}(\tau)$ is the fluctuation of the information on node $j$ at time $t$ induced by a fluctuation of the initial information $\delta s_{i}(0)$ on node $i$, keeping as fixed the initial condition on the other nodes. Hence, the larger $\mathcal{D}_{ij}(\tau)$, the less the communicability on a time $\tau$ between nodes $i$ and $j$. It has to be emphasized that $\mathcal{D}_{ij}(\tau)$ defines an ultrametric distance, as it satisfies the following conditions at all $\tau>0$: (i) $\mathcal{D}_{ij}\geq 0$ with $\mathcal{D}_{ij}=0$ if $i=j$; (ii) $\mathcal{D}_{ij}=\mathcal{D}_{ji}$, and (iii) $\mathcal{D}_{ij}\leq\max\left(\mathcal{D}_{iz},\mathcal{D}_{zj}\right)$ $\forall z\in{\cal V}$ (where ${\cal V}$ is the set of nodes of the network).
@@
Even though complex networks lack an explicit spatial embedding, $\mathcal{\hat{D}}$ allows us to reduce our problem to traditional clustering problems in a metric space [37] allowing for the construction of a dendrogram at arbitrary $\tau$  [^1] In fact, $\mathcal{\hat{D}}$ is a projection of how points (nodes) are distributed in the N-dimensional k-space representation of the network.. To define a stable algorithm for the construction of the dendrogram, we choose the average group clustering algorithm, a compromise between the sensitivity to outliers of the complete clustering and the tendency of single-linkage clustering to inhibit compact clusters [39, 37]. In this way, the merging of two clusters happens when the mean distance $\overline{\mathcal{D}}$ between any node belonging to the first cluster and any node belonging to the second cluster becomes smaller than the chosen threshold $\Delta$. The output of this algorithm will thus be a hierarchical tree for each resolution scale $\tau$, without null-model assumptions.
At this point, three natural questions arise: (1) Given the dendrogram at diffusion time $\tau$, is there an optimal partition or cut? (2) Since the metric $\mathcal{\hat{D}}(\tau)$ may, at different $\tau$, determine different dendrograms with variable composition and length of branches (such as in Markov stability [21]), is there an optimal choice of $\tau$? (3) Is it stable the optimal cut at varying $\tau$? We face now the first two questions, leaving the third one in a section below.
@@
Figure 2(a) shows the hierarchical dendrogram for Zachary’s karate club [2, 40], which serves as a benchmark showing the usual partition of the network in two communities. Figure 2(b) shows the entropic phase transition [33] for Zachary’s karate club using the Laplacian entropy $S(\tau)$ in time (red dashed line) and of the entropic susceptibility (i.e., the specific heat) $C(\tau)$.
@@
We propose a comparative standard for selecting the optimal cut quantifying the quality of the different network partitions encrypted in every dendrogram at a fixed diffusion time $\tau$. In particular, we introduce the Partition Stability Index $\Psi(n;\tau)$, as the ratio between the dendrogram gaps and the total dendrogram length, logarithmically measured. Then,
@@
where $\Delta_{n}(\tau)$ is the value of the threshold $\Delta$ corresponding to the $n^{th}$ dendrogram branching, starting from the network as a single cluster (with this choice of the index $n$, $\Delta_{n}(\tau)$ is a decreasing function of $n$) and $\mathcal{N}=[\log\Delta_{1}(\tau)-\log\Delta_{n_{max}}(\tau)]^{-1}$ is a normalization constant.
@@
The quantity $\Psi(n;\tau)$ measures the minimal distance (dissimilarity) in logarithmic scale between all the pairs of clusters belonging to level $n+1$ and to level $n$. The logarithmic scale is motivated by the definition of the entropy $(S(\tau)\sim\left<\log\hat{\rho}(\tau)\right>_{\tau})$ itself: the higher $\Psi(n;\tau)$, the higher the (Laplacian) information gap between two consecutive branching indices $n$ and $n+1$. It is, therefore, natural to consider the optimal cut of the dendrogram at the value $n$ maximizing $\Psi(n;\tau)$, which contains the maximal Laplacian information induced by the network’s topology. In the case of multiple intrinsic scales of the network, $\Psi(n;\tau)$ can present different local maxima, reflecting characteristic mesoscopic network regions.
@@
Figure 2: LRG multi-scale decomposition. (a) Normalized dendrogram for the Zachary’s karate club with $\tau\lesssim\tau^{\prime}=1/\lambda_{max}$. The Red dashed line reflects the optimal division as stated by the maximum of $\Psi$. Different communities are shaded in different colors. (b) Entropy parameter (red dashed line, $(1-S)$), and specific heat (blue solid line, $C$), versus the temporal resolution parameter of the network, $\tau$. Black dashed line indicates $\tau^{\prime}=1/\lambda_{\max}$. The black dotted line indicates the time of the maximum heat capacity, $(\tau^{*})$. (c) Partition Stability Index $(\Psi)$ versus gap number for $\tau\lesssim\tau^{\prime}$. Note the high $\Psi$ values giving place to the usual division in two and four communities of Zachary’s karate club. Inset: Optimal division into communities of the network.
@@
The interpretation of the dendrogram at different $\tau$’s needs of the recently introduced theory of the LRG [17]. It grounds on the fact that $\hat{L}$ –and specifically its eigenvalues $\lambda$ and eigenvectors |$\ket{\lambda}$– encodes the whole information about all the independent network diffusion modes (see Appendix B and Appendix C). As shown in [33] and [17], these modes collected in the operator $e^{-\tau\hat{L}}$ determine the clusters of nodes that are informationally connected at the time-scale $\tau$. Indeed, the LRG takes advantage of the fact that, varying the diffusion time $\tau$, the operator $\hat{\rho}(\tau)$ acts as a sort of “scanner” of the characteristic network scales [41]. Such mesoscopic scales can be detected by progressively integrating out smaller and smaller Laplacian eigenmodes and, consequently, the corresponding network structures as $\tau$ increases. For finite but large networks, the susceptibility $C(\tau)$, determines the existence of a resolution window where the different partitions of the network in clusters are relevant. The smallest possible network scale corresponds to the maximum eigenvalue $\lambda_{max}$: the choice of a diffusion time-scale $\tau\sim\nicefrac{{1}}{{\lambda_{max}}}$ coincides with the finest possible resolution of the network structures (this sheds light on the previously enigmatic origin of the resolution limit on the community detection field [42]). Starting from this scale $(\tau^{\prime}=1/\lambda_{max})$, one can consider progressively larger diffusion times integrating out smaller network eigenmodes up to $\tau\sim 1/\lambda_{gap}$, where $\lambda_{gap}$ is the smallest positive eigenvalue, also called Fiedler eigenvalue or gap, and monitoring the stability across diffusion scales of the partition of the network into clusters (i.e., defocusing microscopic details). This is related to the maximal characteristic network scale marked by the specific heat peak at long times $(\tau^{*})$, beyond which it rapidly decreases to zero.
@@
Figure 2(c) shows the behavior of $\Psi(n)$ for Zachary’s karate club and the inset the optimal partition for the time $\tau$ indicated with the vertical dashed line in Fig. 2(b). We have also analyzed the structural organization of two paradigmatic hierarchical nested networks: hierarchical-modular networks with a preferential attachment rule (HM-CP, producing a core-periphery structure involving central connector hubs with local and global rich clubs [48]) and the Dorogovtsev-Goltsev-Mendes graph (DGM), a pseudo-fractal graph with high clustering properties [49]. As better illustrated in Appendix E, our LRG-based approach can capture the full intrinsic modular structure of nested hierarchical networks (HM-CP), even when the inter-module connections and communication paths tend to be centralized through the hubs, as observed in real neural and brain networks [50]. The same happens for the highly hierarchical DGM graph, whereas usual community detection methods present severe issues in unraveling static communities being unable to disentagle the complete DGM hierarchical structure (see Appendix E). Figures 3(a) and (b) illustrate the emergent nested structure hinted in the distance matrix, $\mathcal{\hat{D}}$, for HM-CP and DGM networks. Finally, we have tested the Lancichinetti-Fortunato-Radicchi benchmark [51], a network generator with a priori known heterogeneous communities, widely used as a stringent criterion to compare different community detection algorithms (see Appendix E).